OpenStudy (anonymous):

36x^2-12x+1=4 What Are the solutions of this equation?????? Please help

OpenStudy (anonymous):

\[36x^2-12x-3=0\]

OpenStudy (anonymous):

Can u help me with another one plz.

OpenStudy (anonymous):

actually we still have to find the answer

OpenStudy (anonymous):

i know

OpenStudy (anonymous):

when were finished

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so what do i do?

OpenStudy (anonymous):

use the quadratic equation formula

OpenStudy (anonymous):

wow wtf u took that long to tell me that.

OpenStudy (anonymous):

no..I was just doing some math for easy calc

OpenStudy (anonymous):

sorry any way..

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

Which value completes the square for x^2-3x

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

how would u solve

OpenStudy (anonymous):

I will try..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

got it..:D

OpenStudy (anonymous):

let us assume that it is (x-K)^2

OpenStudy (anonymous):

(x-K)^2=x^2-2xK+K^2

OpenStudy (anonymous):

our polynomial goes like x^2-3x

OpenStudy (anonymous):

so let 2xK=3x --->K=3/2

OpenStudy (anonymous):

\[(x-3/2)^2=x^2-(2*(3/2)*x)+9/4)\]

OpenStudy (anonymous):

so the number is 9/4

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

My pleasure:D